# PixelWorld: Towards Perceiving Everything as Pixels

### v2

@@ 6 Conclusion
We present PixelWorld, a benchmark that renders text, tables, code, and images as pixels, enabling direct evaluation of the Perceive Everything as Pixels (PEAP) paradigm. Experiments yield three takeaways: (1) Semantic understanding. PEAP matches token baselines on sentence-/paragraph-level tasks, while its patch-level attention closely mirrors token attention, pointing toward “vision-as-token” models. (2) Reasoning. Accuracy drops on math, logic, and program-repair benchmarks; Chain-of-Thought prompts narrow but do not eliminate this gap. (3) Multimodal tasks. Pixel input outperforms OCR pipelines on websites, slides, and documents by preserving spatial context and avoiding recognition errors. To curb the higher latency of pixel inputs, we introduce PEAP-Fast, which prunes blank patches and accelerates inference by up to $3\times$3 × without hurting accuracy. Taken together, these findings underscore both the promise and trade-offs of PEAP. PixelWorld thus serves as a practical stress test and diagnostic benchmark, encouraging the community to adopt PEAP as a holistic yardstick while guiding research on efficiency improvements and on closing the reasoning gap in next-generation multimodal agents.
Limitations
Due to constraints in time and computational resources, we could not comprehensively evaluate all benchmarks and language models. To simplify our experimental setup, we primarily focused on tasks that can be evaluated through straightforward string-based computations. Tasks involving complex settings such as function calling, code reasoning, or retrieval remain insufficiently covered.
Furthermore, unless prompts were explicitly provided by the dataset, we employed uniform Direct Prompting and Chain-of-Thought prompting strategies, which may not be representative enough and could introduce biases. More sophisticated prompting techniques, such as ReAct, were not thoroughly evaluated.
Finally, our framework focuses specifically on evaluating the textual reasoning abilities of vision-language models, leaving other multimodal models—such as video generation models that lack textual reasoning capability—beyond the scope of this evaluation.
Ethics Statement
This paper focuses on the reasoning capabilities and interpretability of vision-language models. Vision-language models possess strong abilities for understanding and interacting with the real world, and thus carry potential risks of misuse. However, our methods concentrate on benign reasoning tasks and propose interpretability mechanisms that can help mitigate potential misuse.
Additionally, the datasets used in our experiments primarily comprise English data. While our framework is potentially transferable to other languages, vision-language models may suffer from reduced performance due to limited availability of resources in languages other than English.
Reproducibility Statement
All code used in this paper has been made publicly available, along with intermediate results of all synthetic data. Additionally, we set the inference temperature uniformly to 0 across all models. Thus, we believe our work is highly reproducible. However, minor discrepancies may arise due to differences in hardware and inference frameworks, potentially leading to results that are not strictly identical.
@@ Appendix B Attention Heatmap before and after ImageFast Method
Appendix C Model Size and Computational Budget
All data syntactic process and experiments used around 50 hours of a 32-core CPU host and around 200 GPU hours on Nvidia A6000. The total cost of API calls was less than 1,000 USD.
Appendix D Usage of AI Assistants
During the development process, we drafted the code framework and data structures by ourselves initially, using Copilot to assist in enhancing development efficiency. Additionally, in the writing phase, we only employed ChatGPT for proofreading purposes, such as correcting grammatical errors and properly formatting tables.

### v3

@@ 6 Conclusion
We present PixelWorld, a benchmark that renders text, tables, code, and images as pixels, enabling direct evaluation of the Perceive Everything as Pixels (PEAP) paradigm. Experiments yield three main takeaways. (1) Semantic understanding: PEAP performs on par with token-based baselines on sentence- and paragraph-level tasks, and its patch-level attention exhibits similar global structures to token attention, suggesting partial transfer of language modeling behavior into the visual domain. (2) Reasoning: Performance drops on math, logic, and program-repair benchmarks, though Chain-of-Thought prompting mitigates but does not close this gap, highlighting the continued need for explicit reasoning structure. (3) Multimodal perception: Pixel-based inputs outperform OCR pipelines on websites, slides, and documents by preserving spatial context and avoiding recognition noise. To alleviate the higher computational cost of pixel inputs, we propose PEAP-Fast, which prunes blank patches and achieves up to a $3\times$ inference speedup with minimal accuracy loss. Overall, these findings illustrate both the potential and limitations of PEAP, positioning PixelWorld as a diagnostic and reproducible benchmark for studying unified vision–language representations and guiding future work on efficiency and reasoning in multimodal agents.
@@ Appendix B Attention Heatmap before and after ImageFast Method
Appendix C Attention Heatmap Comparison Between Datasets
We provide additional attention visualizations on three representative datasets—MBPP, MMLU-Pro, and MathVerse—to illustrate how attention patterns vary across program synthesis, reasoning, and STEM tasks.
Figure 11: Example 1: Last Layer Attention Heatmap on Qwen2VL-7B from MBPP.
Figure 12: Example 2: Last Layer Attention Heatmap on Qwen2VL-7B from MMLU-Pro.
Figure 13: Example 3: Last Layer Attention Heatmap on Qwen2VL-7B from MathVerse.
Appendix D Broader Impact Statement
This work explores a unified pixel-based perception paradigm that eliminates the need for separate text and image tokenization. While such unification can simplify multimodal pipelines and reduce reliance on OCR systems, it also introduces new risks. The computational cost of pixel-based models remains significantly higher than text-based counterparts, which may limit accessibility and increase the carbon footprint of large-scale training and deployment. Furthermore, because pixel inputs may contain sensitive visual information, researchers must ensure that data synthesis and collection comply with privacy and ethical standards.
On the positive side, the PixelWorld benchmark provides a transparent and reproducible foundation for assessing multimodal understanding, encouraging fair comparisons across models and modalities. By highlighting where pixel-based representations succeed and fail, this work aims to guide the community toward more efficient and interpretable multimodal systems, fostering broader exploration of unified perception without compromising ethical or environmental considerations.